In the spin cycle of a washing machine, the tub of radius 0.44 m rotates at a constant rate of 705 rev/min. What is the centrifugal forcess acts on clothes having capacity 7 Kg.Answer in units of N.
1
Expert's answer
2011-12-01T08:50:30-0500
Let's make some denotations:

R = 0.44 m;
M = 7 kg;
N = 705 rev/min.

The liniar velocity of a point at the tub is
V = N*2*R*pi = 705*2*0.44*3.14 = 1948.056 m/min = 32.4676 m/s.
Centripetal acceleration:
Ac = V²/R = 32.4676²/0.44 = 2395.784204 m/s².
So, centrifugal force is
F = M*Ac = 7*2395.784204 = 16770.489428 N.